OpenStudy (anonymous):

multiply both sides of an equation by 0 to solve it.

OpenStudy (texaschic101):

what ?

OpenStudy (anonymous):

lol...

hartnn (hartnn):

You have great potential to become the world's best mathematician! Let such great ideas keep on flowing! :)

OpenStudy (anonymous):

@hartnn ... OS has got alot of talent....:D

OpenStudy (anonymous):

certainly you are genius.

OpenStudy (anonymous):

@lilmac10423 we can't multiply any equation with 0.. :p

Parth (parthkohli):

(hint: the post is a joke.)

OpenStudy (anonymous):

good joke

Parth (parthkohli):

you're welcome.

OpenStudy (unklerhaukus):

5 = 8 5*0 = 8*0 0 = 0 True it works!

OpenStudy (anonymous):

@ParthKohli ... that made the question easy...:d ty

OpenStudy (anonymous):

u'll get a big o by the examiner @UnkleRhaukus easy 4 examier

OpenStudy (anonymous):

ok bye

OpenStudy (unklerhaukus):

a big o is better than a little o

OpenStudy (anonymous):

yeah eggs

OpenStudy (anonymous):

4 fun

Parth (parthkohli):

@UnkleRhaukus Please convince my parents. I've got a big 0 on my report card.

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

say them.. that someone multiplied ur marks with 0 :p using the Lilmac's postulates :p

OpenStudy (anonymous):

good one @Shizen

OpenStudy (unklerhaukus):

make it into a smiley